OpenStudy (anonymous):

Which of the following best describes the solution to the equation below? 30x + 27 = -45x - 6

OpenStudy (redheadangel):

First you have to move -45 to the other side of the equal side

OpenStudy (anonymous):

ok

OpenStudy (redheadangel):

you would do that by adding 45 to 30x

OpenStudy (redheadangel):

75x+27=-6

OpenStudy (redheadangel):

then you subtract 27 from -6

OpenStudy (anonymous):

Ok. im writing this down

OpenStudy (redheadangel):

that would get you 21 so...75x=21

OpenStudy (redheadangel):

then divide 75 from 21

OpenStudy (redheadangel):

that would mean your answer would be: 3.57

OpenStudy (anonymous):

these are the choices exactly one real solution, exactly one real solution, no real solutions infinite real solutions

OpenStudy (redheadangel):

oh.... ok

OpenStudy (redheadangel):

so: exactly one real solution

OpenStudy (anonymous):

The first one is x = -11/25 The second one is x = 2 1/5

OpenStudy (redheadangel):

?

OpenStudy (anonymous):

The first Choice includes the 1st one and the second choice included the other one

OpenStudy (redheadangel):

ohh.. hmmm

OpenStudy (anonymous):

30x + 27 = -45x - 6 30x + 45x = -6 -27 75x = 33 x = 33/75

OpenStudy (redheadangel):

yes, but that isn't one of the answers..

OpenStudy (anonymous):

exactly

OpenStudy (redheadangel):

I am not sure where to go from here maybe @Here_to_Help15 can take over :)

OpenStudy (anonymous):

hold on x = 33/75 x = 11/25

OpenStudy (anonymous):

exactly one real solution

OpenStudy (redheadangel):

ohhhh so the first one then @Kyliejones

OpenStudy (anonymous):

Ok thanks!

OpenStudy (here_to_help15):

30x + 27 = -45x - 6 30x + 45x = 27 - 6

OpenStudy (here_to_help15):

30x + 27 = -45x - 6 * 30x + 45x = 27 - 6 i meant hehe

OpenStudy (here_to_help15):

?

OpenStudy (here_to_help15):

75x=21 so x=21/75 can you go on